# Addressing Positivity Violations in Continuous Interventions through Data-Adaptive Strategies

### v2

@@ Appendix D Simulation 1: Bootstrap CI coverage probabilities
Figure 8: Bootstrap CI coverage probabilities in Simulation 1. Empirical coverage of 95% bootstrap confidence intervals for three estimators—Standard (orange), Feasible (red), and Trimming (green)—as a function of the intervention value AA. Each panel corresponds to a different true outcome function: Linear, Logarithmic, or Sine. The horizontal dashed line indicates the nominal 95% coverage level. When the outcome is linear (correctly specified by the model), all estimators achieve near-nominal coverage. Under nonlinear outcomes (model misspecification), the Standard estimator shows undercoverage in regions with limited support, while the Trimming estimator attains nominal coverage in well-supported areas but becomes unstable near the boundaries. The Feasible estimator provides stable coverage across the intervention range, achieving nominal levels when the outcome model is correctly specified, and offering more reliable coverage under misspecification.

### v3

@@ Appendix G Simulation 1: Bootstrap CI coverage probabilities
Figure 8: Bootstrap CI coverage probabilities in Simulation 1. Empirical coverage of nominal 95% nonparametric bootstrap confidence intervals for three estimated curves—the standard CDRC plug-in estimator $\hat{m}_{n}^{\mathrm{standard}}$ (orange), the feasible curve estimator $\hat{m}^{\mathrm{feasible}}_{n,\alpha}$ (red), and the trimmed curve estimator $\hat{m}^{\mathrm{trim}}_{n,\alpha}$ (green)—evaluated pointwise over intervention values AA. Coverage is computed with respect to the corresponding true target curves under the data-generating mechanism (i.e., the CDRC, feasible curve, and trimmed curve, respectively). Each panel corresponds to a different outcome regression in the data-generating process (Linear, Logarithmic, Sine). The horizontal dashed line marks the nominal 95% level. When the outcome model is linear (and thus correctly specified by the working parametric outcome regression), coverage is close to nominal for all three estimators. Under nonlinear outcomes (model misspecification), coverage of the standard CDRC estimator deteriorates in regions with limited practical overlap, while the trimmed estimator attains near-nominal coverage in well-supported regions but becomes unstable near the boundaries. The feasible estimator remains comparatively stable across the intervention range and exhibits more reliable pointwise coverage under misspecification.